OpenStudy (anonymous):

Calculus lim as x approaches infinity x^2-5/4 I need help!!!

OpenStudy (lgbasallote):

\[\lim_{x\rightarrow \infty} \frac{x^2 - 5}4\] ???

OpenStudy (anonymous):

yes i need help :(

OpenStudy (lgbasallote):

don't you just substitute infinity into x?

OpenStudy (anonymous):

that's what i thought but then what would my result be is there any other way that you know of?

OpenStudy (anonymous):

the limit is infinite..

OpenStudy (anonymous):

indeed. it is easier to see if you split it up and say lim to infinity of (x^2/4)-5/4... from there wecan plug in increasing numbers for x to see that the limit is infinity...

OpenStudy (lgbasallote):

well there is another way...but it results the same....

OpenStudy (lgbasallote):

\[\lim_{x\rightarrow \infty} \frac{x^2 - 5}4\] divide everything by x^2 \[\Large \implies \lim_{x\rightarrow \infty} \frac{1 - \frac 5{x^2}}{\frac 4{x^2}}\] if you put in infinity...same result...

OpenStudy (anonymous):

^true igbassolat is right. that is the best way to do these infinity limits.. because not all functions will be as easy to see the limit right away...

OpenStudy (anonymous):

hmm im sorry but would that be my final answer? why am i dividing by x^2?